# Privacy-Preserving Training-as-a-Service for On-Device Intelligence: Concept, Architectural Scheme, and Open Problems

### v1

@@ V Conclusions and Future Works
In this paper, we propose Privacy-Preserving Training-as-a-Service (PTaaS), a novel service computing paradigm for on-device intelligence (ODI). PTaaS aims to address the challenges faced by on-device model training with privacy guarantees by outsourcing the training of end-side artificial intelligence (AI) models to cloud or edge-based service providers, with only anonymous queries shared with remote servers. PTaaS enables the development of high-performance, customized on-device AI models while ensuring data privacy and alleviating the computation, storage, and energy constraints of end devices. Future works include enhancing privacy protection mechanisms, optimizing cloud-edge resource collaborative management, improving customized model training, and establishing standard specifications to ensure sustainable development of PTaaS.
@@ Acknowledgement
We thank GPT-4 and Claude-3 for their invaluable contributions as brainstorming partners and their assistance in language refinement.

### v2

@@ V Conclusions and Future Works
In this paper, we propose Privacy-Enhanced Training-as-a-Service (PTaaS), a novel service computing paradigm for on-device intelligence (ODI). PTaaS aims to address the challenges faced by on-device model training with enhanced privacy by outsourcing the training of end-side artificial intelligence (AI) models to cloud or edge-based service providers, with only anonymous queries shared with remote servers. PTaaS enables the development of high-performance, customized on-device AI models with enhanced data privacy while alleviating the computation, storage, and energy constraints of end devices. Future works will focus on enhancing privacy protection mechanisms, optimizing cloud-edge resource collaborative management, improving customized model training, designing novel pricing strategies, and establishing standard specifications, so as to ensure the sustainable development of PTaaS.
@@ Acknowledgement
We thank GPT-4 and Claude-3 for their invaluable contributions as brainstorming partners and their assistance in language refinement. We also thank professor Zhuotao Liu from Tsinghua University for inspiring suggestions.